# ProxNF: Neural Field Proximal Training for High-Resolution 4D Dynamic Image Reconstruction

### v1

@@ Abstract
Accurate spatiotemporal image reconstruction methods are needed for a wide range of biomedical research areas but face challenges due to data incompleteness and computational burden. Data incompleteness arises from the undersampling often required to increase frame rates and reduce acquisition times, while computational burden emerges due to the memory footprint of high-resolution images with three spatial dimensions and extended time horizons. Neural fields, an emerging class of neural networks that act as continuous representations of spatiotemporal objects, have previously been introduced to solve these dynamic imaging problems by reframing image reconstruction to a problem of estimating network parameters. Neural fields can address the twin challenges of data incompleteness and computational burden by exploiting underlying redundancies in these spatiotemporal objects. This work proposes ProxNF, a novel neural field training approach for spatiotemporal image reconstruction leveraging proximal splitting methods to separate computations involving the imaging operator from updates of the network parameter. Specifically, ProxNF evaluates the (subsampled) gradient of the data-fidelity term in the image domain and uses a fully supervised learning approach to update the neural field parameters. By reducing the memory footprint and the computational cost of evaluating the imaging operator, the proposed ProxNF approach allows for reconstructing large, high-resolution spatiotemporal images. This method is demonstrated in two numerical studies involving virtual dynamic contrast-enhanced photoacoustic computed tomography imaging of an anatomically realistic dynamic numerical mouse phantom and a two-compartment model of tumor perfusion.
@@
Dynamic imaging, Neural fields, Proximal method, Stochastic optimization, Spherical Radon transform, Computer-simulation study
@@ I Introduction
Dynamic image reconstruction with high spatial and temporal resolution is important for understanding the evolution of many biological systems [1, 2, 3]. In many dynamic image reconstruction problems, measurements are undersampled and frame-by-frame reconstruction methods yield unsatisfactory resolution. In such cases, higher resolution can be achieved by utilizing a spatiotemporal reconstruction method in which the image is treated as a single spatiotemporal object [4]. However, large-scale spatiotemporal image reconstruction can quickly become computationally prohibitive and exceed memory requirements in three spatial dimensions. Therefore, high-resolution dynamic image reconstruction needs to exploit underlying redundancies in the image to reduce the number of parameters necessary for storing the image [5, 6].
@@
On the one hand, optimization-based image reconstruction that employ semiseparable or low rank approximations has often been advocated to address these challenges. Previous work has demonstrated the use of such approaches to enable several dynamic imaging applications, including X-ray tomography [7], positron electron tomography [8], magnetic resonance imaging[9, 6], and more recently also photoacoustic computed tomography [10, 11, 12]. On the other hand, the use machine learning, and neural fields in particular, has been actively investigated and shows great promise for the solution of this class of image reconstruction problems. Neural fields (NFs) are a particular class of neural networks that provide a continuous representation of an object in space and time. Using a neural field reduces a dynamic image reconstruction problem into estimating network weights by solving a nonconvex optimization problem [13, 14, 15]. Unlike other deep learning methods, NFs are self-supervised. The neural field model is trained only using the measurement data and does not require ground truth images, or labels[16]. Additionally, NFs allow for memory-efficient representation of the object [17, 18] and provide implicit regularization based on the neural architecture[19].
@@
Previous works with NFs utilized first-order (accelerated) gradient based methods [18, 19, 20, 21, 22, 17, 23], such as Adam [24], to train the network parameters. However, training a NF for image reconstruction in this fashion requires the imaging operator and its adjoint to be computed at each gradient update, greatly slowing down training [13, 25, 14]. Furthermore, while computational efficient implementation of discrete-to-discrete imaging operators are readily available, the implementation of a continuous-to-discrete imaging operator acting on a NF representation incur severe computational and memory burden due to the need of repeated evaluation of the NF [13].
@@
This work proposes ProxNF, a proximal splitting method for self-supervised training of a NF to enable memory-efficient, high-resolution spatiotemporal image reconstruction. The key contribution of this splitting method is a separation of concerns between the updates of NF parameters and calls to the imaging operator. In this approach, an update direction in the image domain is computed using the gradient of the data-fidelity term, which requires an evaluation of the discrete-to-discrete imaging operator and its adjoint. Once this update direction is formed, no further computations involving the imaging operator are needed. The NF parameters are then updated independently from the imaging operator by solving a supervised subproblem (the proximal operator). In essence, this method reframes the original self-supervised problem to solving a series of supervised subproblems while reducing the number of calls to the imaging operator and thus the computational burden. Furthermore, the update direction can be computed in a memory efficient manner with framewise-batched gradients of the data-fidelity term and imaging operator. ProxNF allows for reconstruction of large, high-resolution spatiotemporal images that were previously intractable. This method is demonstrated with dynamic image reconstructions in two and three spatial dimensions inspired by dynamic contrast-enhanced photoacoustic tomography imaging of tumor perfusion.
@@
The remainder of this paper is structured as follows. In Section II, spatiotemporal image reconstruction is formulated as an optimization problem and reconstruction methods utilizing NFs for representing the dynamic object are introduced. In Section III, the proposed ProxNF method for training a NF for spatiotemporal image reconstruction is derived. This section also discusses a specific NF architecture, the partition of unity network, and introduces an adaptation to this architecture for representing high-resolution, high-dimensional dynamic images. In Section IV, the design of two numerical studies is presented along with the computational framework for virtual dynamic contrast-enhanced photoacoustic imaging of a numerical tumor perfusion phantom in two and three spatial dimension. The first study assesses the representation power of the proposed NF architecture applied to the dynamic tumor perfusion numerical phantom compared to a class of semiseparable approximations and to the original partition of unity network architecture. The second study compares dynamic image reconstruction in two and three spatial dimensions using the proposed NF representation with ProxNF and a model-based reconstruction method using nuclear norm regularization. In Section V, the results of these numerical studies are presented. Section VI presents the conclusions drawn from these results and future extensions.
@@ II-A1 Continuous-to-discrete formulation
Image reconstruction aims to estimate an object $f$ based on a set of indirect measurements $\boldsymbol{d}$. The object $f$ belongs to an infinite dimensional Hilbert space $\mathbb{U}$. The measurements $\boldsymbol{d}$ belong to a finite dimensional space $\mathcal{D}$ and are associated to $f$ via a possibly nonlinear imaging operator $\mathcal{H}:\mathbb{U}\rightarrow\mathcal{D}$, also known as a forward operator, via
@@
This works is concerned with the specific case of dynamic image reconstruction. That is, $f$ is a spatiotemporal function belonging to $\mathbb{U}=\mathbb{L}^{2}(\Omega_{T})$ with $\Omega_{T}=\Omega\times[0,T]$, where $\Omega\subset\mathbb{R}^{d}$ ($d=2$ or 3) denotes the spatial domain and $T$ is the time horizon.
@@
where $t_{k}=(k-1)\Delta T\in[0,T]$ is the sampling time of frame $k$, $\Delta T$ is the sampling interval, $f(\cdot,t_{k})$ is the object at frame $k$, and $\boldsymbol{n}_{k}$ represents measurement noise. For several applications of practical relevance the measurement aperture translates or rotates between imaging frames [4, 26, 27, 28, 12, 11], yielding imaging operators $\{\mathcal{H}_{k}\}_{k=1}^{K}$ that are different for each frame and measurements that are severely undersampled. In what follows, the notations
$\begin{array}[]{l}\displaystyle\boldsymbol{d}=\sum_{k=1}^{K}\boldsymbol{d}_{k}\otimes\mathbf{e}_{k},\quad\boldsymbol{n}=\sum_{k=1}^{K}\boldsymbol{n}_{k}\otimes\mathbf{e}_{k},\\ \displaystyle\mathcal{H}(f)=\sum_{k=1}^{K}\mathcal{H}_{k}(f(\cdot,t_{k}))\otimes\mathbf{e}_{k}\end{array}$ (3)
are introduced to write the dynamic imaging model in Eqn. (2) in the general form of Eqn. (1). Above, $\{\mathbf{e}_{k}\}_{k=1}^{K}$ indicates the canonical basis of $\mathbb{R}^{K}$ and the symbol $\otimes$ represents the outer product of two vectors yielding a rank-1 matrix.
@@
Above, the cost functional $\mathcal{J}:\mathbb{U}\rightarrow\mathbb{R}$ is the sum of the data-fidelity functional $\mathcal{L}:\mathbb{U}\rightarrow\mathbb{R}$, which is minimized when $\mathcal{H}(f)=\boldsymbol{d}$, and the regularization functional $\mathcal{R}:\mathbb{U}\rightarrow\mathbb{R}$, which encodes prior information on the sought-after image. For the case of Gaussian additive noise, the data-fidelity functional $\mathcal{L}(f)=\sum_{k=1}^{K}\mathcal{L}_{k}(f(\cdot,t_{k}))$ is given by
$\begin{array}[]{cl}\displaystyle\mathcal{L}(f)&=\frac{1}{2\sigma^{2}}\|\mathcal{H}(f)-\boldsymbol{d}\|^{2}\\ &=\sum_{k=1}^{K}\frac{1}{2\sigma^{2}}\|\mathcal{H}_{k}(f(\cdot,t_{k}))-\boldsymbol{d}_{k}\|^{2},\end{array}$
where the weight $\sigma^{2}$ represents the marginal variance of the additive noise $\boldsymbol{n}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})$.
@@ II-A2 Discrete-to-discrete formulation
To derive the discrete-to-discrete formulation, the spatial domain $\Omega$ is discretized using a Cartesian grid with $M$ voxels. The coordinates of the voxel centers are denoted with $\{\boldsymbol{r}_{m}\}_{m=1}^{M}$. The spatial and temporal expansion functions $\beta_{m}(\boldsymbol{r})$ and $\gamma_{k}(r)$ are chosen as the indicator functions of the $m$-th spatial voxel ($m=1,\ldots,M$) and the $\left(t_{k}-\frac{1}{2}\Delta T,t_{k}+\frac{1}{2}\Delta T\right)$ time interval ($k=1,\ldots,K$), respectively. The finite dimensional subspace $\mathbb{U}_{M,K}\subset\mathbb{U}$ consists of all functions $f_{M,K}$ of the form
@@
where ${\rm f}_{m,k}\in\mathbb{R}$ ($m=1,\ldots,M$, $k=1,\ldots,K$) are the expansion coefficients. Eqn. (5) defines an isomorphism $\mathcal{I}:\mathbb{R}^{M\times K}\mapsto\mathbb{U}_{M,K}$ mapping the matrix $\mathbf{F}$ with entries $\{{\rm f}_{m,k}\}_{m,k}$ to the function $f_{M,K}(\boldsymbol{r},t)$. The inner product in $\mathbb{U}_{M,K}$ is defined as
@@
$\Pi_{M,K}^{*}(\mathbf{F})=\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{1}{V_{m,k}}{\rm f}_{m,k}\beta_{m}(\boldsymbol{r})\gamma_{k}(t).$ (7)
The notations $\mathbf{f}_{k}$ and $\Pi_{M,K}^{k}(f)$ are also introduced to denote $k$-th column of $\mathbf{F}$ and $\Pi_{M,K}(f)$, respectively.
@@
$\mathbf{H}_{k}=\sum_{m=1}^{M}\mathcal{H}_{k}(\beta_{m}(\cdot))\otimes\mathbf{i}_{m},$ (8)
where $\{\mathbf{i}_{m}\}_{m=1}^{M}$ denotes the canonical basis of $\mathbb{R}^{M}$. The discrete-to-discrete spatiotemporal imaging operator $\mathbf{H}:\mathbb{R}^{M\times K}\mapsto\mathcal{D}$ is defined in a similar manner to its continuous-to-discrete counterpart, c.f. Eqn. (3). Its action on a matrix $\mathbf{F}$ representing the discretized spatiotemporal object function $f$ reads
@@
Using the above notation, the discrete counterpart of the optimization-based formulation of spatiotemporal image reconstruction reads

### v3

@@ Abstract
Accurate spatiotemporal image reconstruction methods are needed for a wide range of biomedical research areas but face challenges due to data incompleteness and computational burden. Data incompleteness arises from the undersampling often required to increase frame rates, while computational burden emerges due to the memory footprint of high-resolution images with three spatial dimensions and extended time horizons. Neural fields (NFs), an emerging class of neural networks that act as continuous representations of spatiotemporal objects, have previously been introduced to solve these dynamic imaging problems by reframing image reconstruction as a problem of estimating network parameters. Neural fields can address the twin challenges of data incompleteness and computational burden by exploiting underlying redundancies in these spatiotemporal objects. This work proposes ProxNF, a novel neural field training approach for spatiotemporal image reconstruction leveraging proximal splitting methods to separate computations involving the imaging operator from updates of the network parameters. Specifically, ProxNF evaluates the (subsampled) gradient of the data-fidelity term in the image domain and uses a fully supervised learning approach to update the neural field parameters. This method is demonstrated in two numerical phantom studies and an in-vivo application to tumor perfusion imaging in small animal models using dynamic contrast-enhanced photoacoustic computed tomography (DCE PACT).
@@
Dynamic imaging, Neural fields, Proximal methods, Stochastic optimization, Photoacoustic tomography, Computer-simulation studies, Tumor perfusion, In-vivo imaging
@@ I Introduction
Dynamic image reconstruction with high spatial and temporal resolution is important for understanding the evolution of many biological systems [1, 2, 3]. In many dynamic image reconstruction problems, measurements are undersampled and frame-by-frame reconstruction methods yield unsatisfactory accuracy. In such cases, higher accuracy can be achieved by utilizing a spatiotemporal reconstruction method in which the image is treated as a single spatiotemporal object [4]. However, large-scale spatiotemporal image reconstruction can quickly become computationally prohibitive and exceed memory requirements in three spatial dimensions. Therefore, high-resolution dynamic image reconstruction needs to exploit underlying redundancies in the image to reduce the number of parameters necessary to represent the image [5, 6].
@@
On the one hand, optimization-based image reconstruction that employs semiseparable or low rank approximations has often been advocated to address these challenges. Previous work has demonstrated the use of such approaches to enable several dynamic imaging applications, including X-ray tomography [7], positron electron tomography [8], magnetic resonance imaging[9, 6], and more recently also photoacoustic computed tomography [10, 11, 12]. On the other hand, the use of machine learning, and neural fields in particular, has been actively investigated and shows great promise for the solution of this class of image reconstruction problems. Neural fields (NFs) are a particular class of neural networks that provide a continuous representation of an object in space and time. Using NFs reduces a dynamic image reconstruction problem into estimating network weights by solving a nonconvex optimization problem [13, 14, 15]. Unlike other deep learning methods, NFs are self-supervised. The NF model is trained only using the measurement data and does not require ground truth images, or labels[16]. Additionally, NFs allow for memory-efficient representation of the object [17, 18] and provide implicit regularization based on the neural architecture[19].
@@
Previous works with NFs utilized first-order (accelerated) gradient-based methods [18, 19, 20, 21, 22, 17, 23], such as Adam [24], to train the network parameters. However, training an NF for image reconstruction in this fashion requires the imaging operator and its adjoint to be computed at each gradient update, greatly slowing down training [13, 25, 14]. Furthermore, while computationally efficient implementations of discrete-to-discrete imaging operators are sometimes available, the implementation of a continuous-to-discrete imaging operator acting on an NF representation incurs severe computational and memory burden due to the need for repeated evaluation of the NF [13].
@@
This work proposes ProxNF, a proximal splitting method for self-supervised training of an NF to enable memory-efficient, high-resolution spatiotemporal image reconstruction. The key contribution of this splitting method is a separation of concerns between the updates of NF parameters and calls to the imaging operator. In this approach, an update direction in the image domain is computed using the gradient of the data-fidelity term, which requires an evaluation of the discrete-to-discrete imaging operator and its adjoint. Once this update direction is formed, no further computations involving the imaging operator are needed. The NF parameters are then updated independently from the imaging operator by solving a supervised subproblem (the proximal operator). In essence, this method reframes the original self-supervised problem as a sequence of supervised subproblems while reducing the number of calls to the imaging operator and thus the computational burden. Furthermore, the update direction can be computed in a memory-efficient manner via randomized evaluation of the data-fidelity term by stochastically subsampling the imaging frames at each iteration. ProxNF allows for reconstruction of large, high-resolution spatiotemporal images that were previously intractable. This method is demonstrated by use of numerical phantom studies in two and three spatial dimensions and an in-vivo imaging experiment of tumor perfusion using dynamic contrast-enhanced photoacoustic tomography (DCE PACT).
Furthermore, the general idea behind the proposed method, i.e. splitting computation involving the imaging operator from the training of the neural network representing the object, is readily extendable to other classes of neural networks, such as deep image priors (DIP, [26]). In the field of medical imaging, in particular, DIPs have been applied to x-ray computed tomography (CT) [27], positron electron tomography (PET) [28], and dynamic magnetic resonance imaging [29]. The DIP approach has several similarities to NFs. These include 1) the use of a neural network to represent the sought-after object, 2) the implicit regularization provided by such representation, and 3) the self-supervised approach to train the network parameters for a specific input (measurement data) by the solution of an inverse problem. The main difference between NFs and DIPs lies in the choice of architecture. While NFs use a (often dense) neural network architecture mapping spatiotemporal coordinates to image intensities at a specific location, DIPs use a deep convolutional neural network to map a random but fixed code vector to the sought-after image/object estimate. Notably, an expectation maximization algorithm attuned for PET imaging was proposed in [28, 30, 31] that leverages the alternating direction method of multiplier (ADMM, [32]) to split computations involving the imaging operator to the training of the neural representation of the sought-after object. Algorithmically, the proposed method differs from previous methods because ProxNF uses a forward-backward splitting approach while [28, 30, 31] employ backward-backward splitting (ADMM). Advantages of the proposed ProxNF are: 1) The use of an NF representation reduces the memory footprint for large-scale high-dimensional images; 2) The introduction of a stochastic approximation of the data fidelity term increases computational efficiency and further reduce memory requirements.
@@
The remainder of this paper is structured as follows. In Section II, spatiotemporal image reconstruction is formulated as an optimization problem and reconstruction methods utilizing NFs for representing the dynamic object are introduced. In Section III, the proposed ProxNF method for spatiotemporal image reconstruction is derived and analyzed. This section also discusses a specific NF architecture, the partition of unity network, and introduces an adaptation to this architecture for representing high-resolution, high-dimensional dynamic images. In Section IV, the design of two numerical studies is presented along with the computational framework for virtual DCE PACT imaging of a numerical tumor perfusion phantom in two and three spatial dimensions. Section V presents the results of these numerical studies. In Section VI, an application to in-vivo DCE PACT imaging of tumor perfusion in small animal models is presented. Section VII presents the conclusions drawn from these results and future extensions.
@@ II-A1 Continuous-to-discrete formulation
Image reconstruction aims to estimate an object $f$ based on a set of indirect measurements $\boldsymbol{d}$. The object $f$ belongs to an infinite-dimensional Hilbert space $\mathbb{U}$. The measurements $\boldsymbol{d}$ belong to a finite-dimensional space $\mathcal{D}$ and are associated to $f$ via a possibly nonlinear imaging operator $\mathcal{H}:\mathbb{U}\rightarrow\mathcal{D}$, also known as a forward operator, via
@@
This work is concerned with the specific case of dynamic image reconstruction. That is, $f$ is a spatiotemporal function belonging to $\mathbb{U}=\mathbb{L}^{2}(\Omega_{T})$ with $\Omega_{T}=\Omega\times[0,T]$, where $\Omega\subset\mathbb{R}^{d}$ ($d=2$ or 3) denotes the spatial domain and $T$ is the time horizon.
@@
where $t_{k}=(k-1)\Delta T\in[0,T]$ is the sampling time of frame $k$, $\Delta T$ is the sampling interval, $f(\cdot,t_{k})$ is the object at frame $k$, and $\boldsymbol{n}_{k}$ represents measurement noise. For several applications of practical relevance, the measurement aperture translates or rotates between imaging frames [4, 33, 34, 35, 12, 11], yielding imaging operators $\{\mathcal{H}_{k}\}_{k=1}^{K}$ that are different for each frame and measurements that are severely undersampled. In what follows, the notation
$\begin{array}[]{l}\displaystyle\boldsymbol{d}\coloneqq\sum_{k=1}^{K}\boldsymbol{d}_{k}\otimes\mathbf{e}_{k},\quad\boldsymbol{n}\coloneqq\sum_{k=1}^{K}\boldsymbol{n}_{k}\otimes\mathbf{e}_{k},\\ \displaystyle\mathcal{H}(f)\coloneqq\sum_{k=1}^{K}\mathcal{H}_{k}(f(\cdot,t_{k}))\otimes\mathbf{e}_{k}\end{array}$ (3)
is introduced to express the dynamic imaging model in Eqn. (2) in the general form of Eqn. (1). Above, $\{\mathbf{e}_{k}\}_{k=1}^{K}$ indicates the canonical basis of $\mathbb{R}^{K}$ and the symbol $\otimes$ represents the outer product of two vectors yielding a rank-1 matrix.
@@
Above, the cost functional $\mathcal{J}:\mathbb{U}\rightarrow\mathbb{R}$ is the sum of the data-fidelity functional $\mathcal{L}:\mathbb{U}\rightarrow\mathbb{R}$, which is minimized when $\mathcal{H}(f)=\boldsymbol{d}$, and the regularization functional $\mathcal{R}:\mathbb{U}\rightarrow\mathbb{R}$, which encodes prior information on the sought-after image. In this work, the data-fidelity functional $\mathcal{L}(f)=\sum_{k=1}^{K}\mathcal{L}_{k}(f(\cdot,t_{k}))$ is chosen as
$\begin{split}\mathcal{L}(f)&\coloneqq\frac{1}{2\sigma^{2}}\|\mathcal{H}(f)-\boldsymbol{d}\|^{2}\\ \displaystyle&=\sum_{k=1}^{K}\frac{1}{2\sigma^{2}}\|\mathcal{H}_{k}(f(\cdot,t_{k}))-\boldsymbol{d}_{k}\|^{2},\end{split}$ (5)
which corresponds to the negative log-likelihood of observing the data $\boldsymbol{d}$ given the object estimate $f$ under the assumption of additive Gaussian measurement noise $\boldsymbol{n}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})$.
@@ II-A2 Discrete-to-discrete formulation
To derive the discrete-to-discrete formulation, the spatial domain $\Omega$ is discretized using a Cartesian grid with $M$ non-overlapping voxels. The coordinates of the voxel centers are denoted by the vectors $\{\boldsymbol{r}_{m}\}_{m=1}^{M}$. The spatial and temporal expansion functions $\beta_{m}(\boldsymbol{r})$ and $\gamma_{k}(r)$ are chosen as the indicator functions of the $m$-th spatial voxel ($m=1,\ldots,M$) and the $\left(t_{k}-\frac{1}{2}\Delta T,t_{k}+\frac{1}{2}\Delta T\right)$ time interval ($k=1,\ldots,K$), respectively. The finite-dimensional subspace $\mathbb{U}_{M,K}\subset\mathbb{U}$ consists of all functions $f_{M,K}$ of the form
@@
where ${\rm f}_{m,k}\in\mathbb{R}$ ($m=1,\ldots,M$, $k=1,\ldots,K$) are the expansion coefficients. Eqn. (6) defines an isomorphism $\mathcal{I}:\mathbb{R}^{M\times K}\mapsto\mathbb{U}_{M,K}$ mapping the matrix $\mathbf{F}$ with entries $\{{\rm f}_{m,k}\}_{m,k}$ to the function $f_{M,K}(\boldsymbol{r},t)$.
The inner product in $\mathbb{U}_{M,K}$ is defined as
@@
$\Pi_{M,K}^{*}(\mathbf{F})\coloneqq\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{1}{V_{m,k}}{\rm f}_{m,k}\beta_{m}(\boldsymbol{r})\gamma_{k}(t).$ (8)
The notations $\mathbf{f}_{k}$ and $\Pi_{M,K}^{k}(f)$ are also introduced to denote $k$-th column of $\mathbf{F}$ and $\Pi_{M,K}(f)$, respectively. The adjoint operator $\left(\Pi_{M,K}^{k}\right)^{*}:\mathbb{R}^{M}\mapsto\mathbb{U}_{M,K}\subset\mathbb{U}$ is then defined as
$\begin{split}\left(\Pi_{M,K}^{k}\right)^{*}(\mathbf{f}_{k})&\coloneqq\sum_{m=1}^{M}\frac{1}{V_{m,k}}{\rm f}_{m,k}\beta_{m}(\boldsymbol{r})\gamma_{k}(t)\\ {}&=\Pi_{M,K}^{*}(\mathbf{f}_{k}\otimes\mathbf{e}_{k}).\end{split}$ (9)
@@
$\mathbf{H}_{k}\coloneqq\sum_{m=1}^{M}\mathcal{H}_{k}(\beta_{m}(\cdot))\otimes\mathbf{i}_{m},$ (10)
where $\{\mathbf{i}_{m}\}_{m=1}^{M}$ denotes the canonical basis of $\mathbb{R}^{M}$. The discrete-to-discrete spatiotemporal imaging operator $\mathbf{H}:\mathbb{R}^{M\times K}\mapsto\mathcal{D}$ is defined in a similar manner to its continuous-to-discrete counterpart, c.f. Eqn. (3). Its action on a matrix $\mathbf{F}$ representing the discretized spatiotemporal object function $f$ is expressed as
@@
Using the above notation, the discrete counterpart of the optimization-based formulation of spatiotemporal image reconstruction is given by